the reasonable and customary fee for anesthesia services for surgery lasting less than 1 hour in a certain region is $425, as specified by an insurer. a particular anesthesiology office is suspected of charging on average more than this amount and a random sample of 45 of their billings show x = $455 and s = $44. can we conclude at the .055 level of significance that this particular anesthesiology office charges on average more than $425
Answer #1

This is the right tailed test .

The null and alternative hypothesis is

H0 :   = 425

Ha : > 425

Test statistic = z

= ( - ) / / n

= (455 - 425) / 44 / 45

Test statistic = 4.57

P(z > 4.57) = 1 - P(z < 4.57) = 1 - 1 = 0

P-value = 0

= 0.055

P-value <

Reject the null hypothesis .

There is sufficient evidence to conclude that the that this particular anesthesiology office charges on average more than $425 .